# Noisy MRI Reconstruction via MAP Estimation with an Implicit Deep-Denoiser Prior

### v1

@@ III-B Training and Implementation
We trained a single noise-adaptive denoiser $f$ (using the LPDSNet architecture [lpds2025] with $\bm{A}=\bm{I}$) on the fully-sampled reconstructions of the T2w brain data with synthetic additive Gaussian white noise (AWGN) in the range $\sigma\in[0,0.5]$. This denoiser was used for both the ImMAP and DDS based inferences for fair comparison. Additionally, we trained separate end-to-end reconstruction networks (LPDSNet [lpds2025]) on noiseless $R=12$ simulated kspace and noisy $R=8$ collected kspace. We obtain estimated coil noise-covariance matrices $\bm{\Sigma}_{y}$ via a wavelet-based multicoil image domain based method, following [Rep2Rep2025]. Hyperparameters of ImMAP were set as shown in Algorithm 1, taking roughly 128 iterations per inference. Hyperparameters of DDS [chung2024decomposed] were set as $M=5$, $\eta=0.8$, iterations (NFE) =99=99 and $\gamma$ set by performing a grid-search around author-reported $\gamma=0.95$, per acceleration rate $R$.
III-C Simulated Noisy Measurement (setting 1)
Fig. 1 shows example results of the proposed ImMAP against the end-to-end LPDSNet [lpds2025] and the diffusion generative model DDS  [chung2024decomposed]. In the high-acceleration noiseless setting (RR=12, $\sigma$=0), the proposed ImMAP recovered finer structural details and reduced blurring compared to LPDSNet, and more accurately reconstructed small structures than DDS. These results align with the natural image literature, confirming that diffusion-based priors improve reconstruction quality under high undersampling.
Fig. 2 shows the evaluation of our method (ImMAP) against the baseline SENSE [espirit] and the diffusion-based DDS under noisy conditions (RR=8, $\sigma$=0.05). SENSE amplified noise, while DDS produced oversmoothed images that lost fine details. In contrast, ImMAP preserved structural sharpness and detail, despite using the same denoiser prior as DDS. Quantitatively, ImMAP achieved an average (NRMSE, SSIM) of (0.1335 0.9603) and DDS an average of (0.1495 0.9489). Furthermore, ImMAP achieves this performance without the hyperparameter tuning and sensitivity of DDS.
[…]
III-D Retrospective Subsampling with Real Noise (setting 2)
Lastly, we evaluated all methods on a test set with $R=8$ retrospective k-space subsampling and real measurement noise. As shown in Figures 3 and 4, SENSE reconstructions exhibit significant noise amplification. LPDSNet yields overly smooth results, consistent with the tendency of end-to-end networks to suppress fine details. While both diffusion-based methods preserve structure without noise amplification, DDS produces more smoothed reconstructions than ImMAP, especially at higher accelerations. Visual inspection confirms that ImMAP recovers more anatomical detail, avoiding both noise and oversmoothing.
[…]
@@ IV Conclusion
We introduced ImMAP, a diffusion-based MRI reconstruction method that integrates acquisition physics directly into a MAP estimation framework. Experiments on both synthetic and real noisy data demonstrate that ImMAP outperforms state-of-the-art deep learning and diffusion-based approaches, preserving anatomical detail without noise amplification or oversmoothing. Notably, it achieves robust performance without the parameter tuning/sensitivity of existing diffusion methods. Together, these results establish ImMAP as a reliable and interpretable baseline, clarifying the practical utility and limitations of diffusion models for accelerated MRI reconstruction.

### v2

@@ III-B Training and Implementation
We trained a single noise-adaptive denoiser $f$ (using the LPDSNet architecture [6] with $\bm{A}=\bm{I}$) on fully-sampled reconstructions of the T1w brain data with synthetic additive Gaussian white noise (AWGN) in the range $\sigma_{t}\in[0,0.5]$. This denoiser was used for both the ImMAP and DDS based inferences for fair comparison. We trained two noise-adaptive end-to-end reconstruction networks (LPDSNet [6]) on $8\times$ accelerated simulated noise data and measurement noise data. Both the denoising LPDSNet (used by ImMAP and DDS) and reconstruction LPDSNet shared the same architecture hyperparameters for fair comparison (30 layers, 169 channels, and $7\times 7$ conv. kernels). Hyperparameters of DDS [1] were set as $M=5$, $\eta=0.8$, 83 iterations, and $\gamma$ set by performing a grid-search per noise-level (see Section III-D).
III-C Joint Reconstruction and Denoising Performance
Fig. 1 shows the quantitative performance of ImMAP against the end-to-end reconstruction network LPDSNet [6] (competitive with E2EVarNet [12]) and the state-of-the-art denoising diffusion model DDS [1]. Performance was measured across three metrics: Normalized Root-Mean-Squared-Error ($\downarrow$] NRMSE), Learned Perceptual Image Patch Similarity (LPIPS) [15] ($\downarrow$] LPIPS), and the Structural Similarity Image Metric [14] ($\uparrow$] SSIM). Across all metrics, ImMAP performs competitively with DDS. ImMAP is favored by LPIPS over DDS, especially in the low-noise range ($\sigma_{y}=0.01$}). The performance on measurement noise the setting ($\sigma_{y}=\text{meas.}$}) heavily favors the end-to-end reconstruction of LPDSNet, likely due to the oversmoothing tendencies of end-to-end networks [1] being favored by the presence of the irreducible noise term in image quality metric computations.
Fig. 2 shows representative visual examples from each reconstruction method across noise-levels, with addition of the classical reconstruction baseline method SENSE [13]. Quantitative metrics are included for each reconstruction. We observe that ImMAP has very strong low-noise ($\sigma_{y}=0.01$}) performance, recovering lesions missed/obscured by SENSE, LPDSNet, and DDS (see image zooms). For the visual on $\sigma_{y}=0.02$}, we observe competitive performance across all quantitative metrics, however, visual inspection reveals a stark difference. LPDSNet is shown to fail to remove aliasing artifacts and DDS is shown to produce noise-like artifacts around the brain edges, while ImMAP achieves superior visual quality. For $\sigma_{y}=\text{meas.}$}, we observe that ImMAP is able to recover fine texture and details while LPDSNet appears to oversmooth image regions and DDS appears again to introduce noise-like artifacts.
III-D Ablation of ImMAP and DDS
The proposed ImMAP demonstrates strong reconstruction stability w.r.t. its noise-injection/stochasticity parameter ($\beta$), where $\beta\rightarrow 0$ achieves maximum noise-injection and $\beta=1$ removes noise-injection. As noise injection is introduced to help escape local-maxima, we expect this parameter to improve reconstruction performance as $\beta\rightarrow 0$. Additionally, $\beta$ implicitly defines a noise-schedule for $\sigma_{t}$, and hence the number of iterations of ImMAP until convergence. This schedule is not followed strictly due to the noise-level estimation of ImMAP (see Line 1 of Alg.1). Fig. 3 plots the quantitative reconstruction performance and iteration count of ImMAP with varying $\beta$. We see that, across noise-levels, performance of ImMAP improves as $\beta\rightarrow 0$, with similar reconstruction performance for a large range of values until $\beta>0.3$. This is with the exception of reconstruction in the measurement noise setting, where irreducible measurement noise is likely favoring a reconstruction with less fidelity at high $\beta$. As $\beta$ increases, we also see that the algorithm takes fewer iterations to converge. As in Kadkhodaie et al. [9], we find that the number of iterations until convergence is much lower than theoretically determined by the $\beta$-defined noise-schedule of $\sigma_{t}$. It is clear that $\beta>0$ should be set as low as possible (0.01 is sufficient), taking into account any possible computational budget.
State-of-the-art diffusion model DDS [1] has its own hyperparameters, of which $\gamma$, corresponding to a Lagrange multiplier term in its conjugate gradient solves, requires careful tuning. Fig. 4 shows the performance of DDS (NRMSE) vs. $\gamma$, across noise-levels. We observe that the optimal point (minimum) of each noise-level curve occurs in a different location. In Section III-C, we used the optimal $\gamma$ determined from these curves. In practice, despite fully-sampled reference data being used for training, fully-sampled reference data at the desired inference noise-level may not be available, making this grid-search for $\gamma$ difficult or uninformative. In contrast, the optimal choice of hyperparameter ($\beta$) for ImMAP is much clearer.
Fig. 5 shows a visual comparison of hyperparameter changes for both DDS and the proposed ImMAP. DDS shows smoothing and artifact introduction with decreasing $\gamma$. Additionally, the optimal setting of $\gamma$ on this example reconstruction is conflicting between image metrics. In contrast, ImMAP shows similar visual quality for both $\beta=0.01$ and $\beta=0.1$, with better quantitative performance across all metrics at lower $\beta$.
@@ IV Conclusion
We introduced ImMAP, a diffusion-based MRI reconstruction method that integrates acquisition physics directly into a MAP estimation framework. Experiments on both simulated noise and measurement noise reconstruction tasks demonstrate that ImMAP outperforms state-of-the-art deep learning and diffusion-based approaches, preserving anatomical detail without noise amplification or oversmoothing. Notably, it achieves robust performance without the parameter tuning/sensitivity of existing diffusion methods. Together, these results establish ImMAP as a reliable and interpretable baseline, clarifying the practical utility and limitations of diffusion models for accelerated MRI reconstruction.